# An Information-Theoretic Regularizer for Lossy Neural Image Compression

### v2

@@ 4.2 Results
In Fig.3, we present the BD-Rate comparison at four training steps: $0.5\times 10^{6}$, $1\times 10^{6}$, $1.5\times 10^{6}$, and $2\times 10^{6}$. The regularization factors $\alpha$ for the hyperprior, autoregressive, and attention models are set to 0.1, 1 and 1, respectively. It can be observed that, at most equal training steps for all three models, our proposed regularization method can yield better compression performance than the vanilla optimization method. In particular, at the $2\times 10^{6}$ step, the BD-Rates for hyperprior, autoregressive, and attention are -0.88%, -0.92% and -0.92%, respectively. For hyperprior and autoregressive, after a brief convergence period, the BD-Rates are gradually improved from 0 to around -1%. For attention, the regularization immediately improves the compression performance at a very early $0.5\times 10^{6}$ training step. In Fig.4, we present the detailed rate-distortion curves at the $2\times 10^{6}$ step on the Kodak dataset. It can be observed that, due to the introduction of the additional conditional source entropy regularization, the bit rate (or the latent entropy) of each point is decreased by a smaller value. From the form of regularization in Eqn.(12), i.e., $-\alpha H(\bm{X}|\bm{\hat{X}})$, the extent of such bit rate reduction depends on both the regularization factor $\alpha$ and the conditional source entropy $H(\bm{X}|\bm{\hat{X}})$. Note that a more advanced conditional source entropy modeling for $q_{\theta}(\bm{X}|\bm{\hat{X}})$ does not experimentally improve the optimization performance. Instead, the alignment with the latent entropy model proves to be more impactful.
Figure 4: Compression performance on Kodak. Our regularizer achieves -1.13%, -1.15%, and -0.84% BD-Rates for the hyperprior, autoregressive, and attention models, respectively.
@@
Test
set
Regularized v.s. Unregularized
hyperprior autoregressive attention
Pixel-style -2.20% -1.96% -0.47%
Screen -1.15% -2.01% -0.69%
Game -1.31% -0.83% -0.82%
Pathology -1.11% -0.81% -2.38%
Natural -0.88% -0.92% -0.92%
@@
In machine learning, regularization plays an essential role in preventing overfitting and enhancing generalization. In Table 1, we directly apply the models trained on natural images (at the $2\times 10^{6}$ step) to the four unseen domains and evaluate the effectiveness of our regularization method. The results indicate that, across all new domains, regularized models consistently outperform unregularized ones. Performance improvements vary by domain and model, with notable gains in some cases. For instance, in the pathology domain with the attention model, the regularization achieves a BD-Rate of -2.38%, more than double the improvement observed in the natural domain.
More experiments.
In Fig.(5), we continue training both the unregularized and regularized hyperprior models to $4\times 10^{6}$ steps and demonstrate the effect of the regularization factor $\alpha$ on the regularization. Herein, with extended training, the regularization consistently enhances optimization. The best performance, with $\alpha=1$, is achieved at the $2.5\times 10^{6}$ step, showing a BD-Rate of -1.24%. As training progresses to $4\times 10^{6}$ steps, the gain reduces to around 0.7%. For the effect of hyperparameter $\alpha$, it is shown that either the value is too large or too small will not benefit the optimization, which can be explained as follows. As discussed in Sec. 3.2, our $\alpha H(\bm{X}|\bm{\hat{X}})$ maximization regularization is essentially a bound upper of the ground-truth objective $\alpha(H(\bm{X}|\bm{\hat{X}})-H(\bm{U}|\bm{\hat{X}}))$. When $\alpha$ is large, this $H(\bm{U}|\bm{\hat{X}})$ gap is also amplified, which harms the regularization accuracy and thus the optimization performance. When $\alpha$ is small, the regularizer has little contribution to the overall loss, and the optimization is reduced to the vanilla objective.
@@
Figure 5: Effects of regularization factor $\alpha$. The compression model is hyperprior and the anchor is training with the vanilla rate-distortion loss (Eqn.(2)) under an equal number of training steps.
Complexity.
Our regularization method is only applied at the network training stage and thus imposes no added inference complexity. As for training, the main overhead of our method comes from training an extra source entropy model $q_{\theta}(\bm{X}|\bm{\hat{X}})$, whose network structure is intentionally aligned with the latent entropy model of CompressAI’s reimplementation in our experiment. For the case of $\lambda=0.0018$, we collect the absolute training time of each model and calculate the ratio. As shown in Table. 2, compared to the unregularized anchor, the training time of our proposed method is increased by 19.5%, 11.3%, and 1.5% for hyperprior, autoregressive, and attention, respectively. Notably, as the compression backbone becomes more complex, the overhead is less prominent.
@@
hyperprior autoregressive attention
Training time 119.5% 111.3% 101.5%
@@ 5 Conclusion
This paper investigates a structural entropy regularizer for lossy neural image compression. Specifically, motivated by the information-theoretic analysis of compression models, we demonstrate that, in addition to minimizing latent entropy, maximizing conditional source entropy can be equally important. Subsequently, by regularizing the neural networks to simultaneously maximize the conditional source entropy during end-to-end optimization, we experimentally observe improvements in both the in-domain compact representation and the out-of-domain generalization abilities of compression models. Additionally, as the primary overhead of this regularizer comes from training an additional source entropy model, the proposed method imposes no added inference complexity, and the training complexity remains negligible when the compression backbone is substantially more complex than the source entropy model, opening up a new space of regularizing neural compression networks.
@@ 6 Proof of Lemma 1
Substituting Eqn.(18) into Eqn.(19), we can conclude
@@ 7 Proof of Lemma 2
Recalling the proof in Sec. 6, since both the analysis transform $\mathrm{T}_{\mathrm{A}}(\cdot)$ and synthesis transform $\mathrm{T}_{\mathrm{S}}(\cdot)$ are deterministic, we still have
@@ 8 Reproduced baselines
We retrain the hyperprior [4], autoregressive [8] and attention [10] models from scratch, adhering to the default implementation and training configurations of CompressAI’s [43]. Four bit-rate points, i.e., $\lambda\in\{0.0018,0.0035,0.0067,0.0130\}$ are trained with $2\times 10^{6}$ steps. The evaluation results are summarized in Table 3. On average, when compared to the pre-trained models of CompressAI’s, our reproduction models yield 0.39%, 0.41%, and -4.11% BD-Rates [46] for the hyperprior, autoregressive, and attention models, respectively. Notably, on the CLIC dataset, our reproduction models outperform their pre-trained counterparts by -1.24%, -1.56%, and -5.68% BD-Rates for the hyperprior, autoregressive, and attention models, respectively. For the attention model, our reproduction model outperforms the pre-trained model on all three evaluation datasets of Kodak, CLIC, and Tecnick by -3.32%, -3.33%, and -5.68% BD-Rates, respectively. The detailed rate-distortion curves for the Kodak dataset are visualized in Fig.6.
9 Conditional source entropy models
Similar to the latent entropy modeling [4, 8, 10], our regularizer involves an additional modeling of the conditional distribution $q_{\theta}(\bm{X}|\bm{\hat{X}})$. Herein, $q_{\theta}(\bm{X}|\bm{\hat{X}})$ follows exactly the same modeling method and the neural network structure as the latent entropy used. In particular, for the hyperprior model, we have
$\text{where}\ (\mu_{1},...,\mu_{N},\sigma_{1}^{2},...,\sigma_{N}^{2})=f_{\theta}(\bm{\hat{X}}),$ (30)
and $f_{\theta}$ is the neural network; $\mathbb{N}$ denotes the Gaussian distribution; $\mu_{i}$ and $\sigma_{i}^{2}$ denote the mean and variance for the $i^{th}$ dimension, respectively; $N$ denotes the dimension of both $\bm{X}$ and $\bm{\hat{X}}$. For both the autoregressive and attention models, the distribution remains Gaussian. However, for the mean and variance, they are determined by the extra spatial context, i.e.,
$(\mu_{1},...,\mu_{N},\sigma_{1}^{2},...,\sigma_{N}^{2})=f_{\theta}(\bm{\hat{X}},\bm{X}_{<i}),$ (31)
where $\bm{X}_{<i}$ denotes the pixels before the current position. Note that in the original attention network [10], the latent is modeled as the mixture of Gaussian with 3 clusters. In CompressAI’s default implementation, this cluster number is simplified to 1. We follow this latent implementation and use an aligned 1-cluster conditional Gaussian distribution for the source modeling. The detailed network structure for $q_{\theta}(\bm{X}|\bm{\hat{X}})$ is depicted in Fig.7, wherein the spatial context is achieved through a $5\times 5$ masked convolution layer (i.e., MaskConv).
Figure 7: Conditional source entropy models for (a) hyperprior; (b) autoregressive and attention; (c) Details of Conv Block 1; (c) Details of Conv Block 2.
10 On the alignment of entropy models
Figure 8: Regularization performance for (a) the hyperprior compression network with the autoregressive regularizer; (b) the autoregressive compression network with the hyperprior regularizer.
In Sec.4, we highlight the benefits of aligned entropy modeling between latent and source distributions. In this section, we extend the analysis to explore cases where the entropy models are unaligned, investigating two scenarios: a compression network equipped with (a) a stronger source entropy model and (b) a weaker source entropy model. In particular, for the entropy models of the hyperprior and autoregressive networks discussed in Sec.9, we consider the autoregressive entropy model with additional spatial contexts to be a more accurate and thus “stronger” model. Thereafter, using the same experimental setting in Sec.4, we train the hyperprior compression network with a autoregressive source regularizer and the autoregressive compression network with a hyperprior source regularizer, and the results are depicted in Fig.8(a) and Fig.8(b), respectively. The values of $\alpha$ are still set to 0.1 and 1 for the hyperprior and autoregressive regularizers, respectively. The performance of the aligned entropy model is also included for reference. From Fig.8, it can be observed that neither a stronger nor a weaker source entropy model consistently facilitates the optimization performance of the compression networks. Specifically, when a stronger autoregressive regularizer is imposed on the hyperprior network, the average BD-Rate is always above zero during the training. When a weaker hyperprior regularizer is employed by the autoregressive network, the optimization is temporarily improved at the $1.5\times 10^{6}$ training step, with a BD-Rate of -0.20%. In contrast, by simply aligning the entropy models, significant performance improvements are observed.

### v3

@@ 4.2 Results
In Fig.4, we present the BD-Rate comparison at four different training steps. The anchor is trained with the vanilla rate-distortion loss (Eqn.(2)) under an equal number of training steps. The regularization factors $\alpha$ for the hyperprior, autoregressive, attention, ELIC, and MLIC++ models are set to 0.1, 1, 1, 0.1, and 0.1, respectively. It can be observed that, for training steps beyond $0.5\times 10^{6}$ across all five models, our proposed regularization method can achieve better compression performance than the vanilla optimization method. In particular, at the $2\times 10^{6}$ step, the average BD-Rates for hyperprior, autoregressive, and attention are -0.88%, -1.06% and -0.92%, respectively; at the $1\times 10^{6}$ step, the average BD-Rates for ELIC and MLIC++ are -1.56% and -3.03%, respectively. Meanwhile, the results demonstrate that the regularization effect varies across models as training progresses: For the hyperprior, autoregressive, and MLIC++ models, after a brief convergence period, the BD-Rates are gradually improved from 0 to around -1%. For the ELIC model, the regularization is most effective during the earlier training stages, achieving over 2.5% BD-Rates saving at the $0.25\times 10^{6}$ step. For the attention model, the regularization achieves a relatively stable coding gain.
In Fig. 5, we present the detailed rate-distortion (RD) curves on the Kodak dataset. For the hyperprior, autoregressive, and attention models, the results correspond to the $2\times 10^{6}$ training step, while for the ELIC and MLIC++ models, the results are from the $1\times 10^{6}$ step. The results indicate that the introduction of the additional conditional source entropy regularization leads to a slight reduction in the bit rate for most of the points, although a few exceptions exhibit an increase due to training noise. From the form of regularization in Eqn.(12), i.e., $-\alpha H(\bm{X}|\bm{\hat{X}})$, the extent of such bit rate reduction depends on both the regularization factor $\alpha$ and the estimated conditional source entropy $H(\bm{X}|\bm{\hat{X}})$.
Figure 5: Compression performance on Kodak. Our regularizer achieves -1.13%, -1.57%, -0.84%, -1.24%, and -1.82% BD-Rates for the hyperprior, autoregressive, attention models, ELIC, and MLIC++, respectively.
@@
Test set Regularized v.s. Unregularized
hyperprior autoregressive attention ELIC MLIC++
Kodak -1.14% -1.57% -0.84% -1.24% -1.82%
Tecnick -0.83% -0.78% -0.91% -2.00% -4.01%
CLIC -0.66% -0.83% -1.01% -1.45% -3.25%
Natural Avg. -0.88% -1.06% -0.92% -1.56% -3.03%
Pixel-style -2.20% -1.96% -0.47% -2.71% -3.08%
Screen -1.15% -2.01% -0.69% -1.02% -1.72%
Game -1.31% -0.83% -0.82% -2.04% -3.44%
Pathology -1.11% -0.81% -2.38% -1.82% -0.43%
@@
In machine learning, regularization plays an essential role in preventing overfitting and enhancing generalization. In Table 1, we directly apply the models trained on natural images to the four unseen domains and evaluate the effectiveness of our regularization method. The results indicate that, across all new domains, regularized models consistently outperform unregularized ones. Performance improvements vary by domain and model, with notable gains in some cases. For instance, in the pathology domain with the attention model, the regularization achieves a BD-Rate of -2.38%, more than double the improvement observed in the natural domain.
@@
Figure 6: (a) Effects of regularization factor $\alpha$. (b) Regularization performance for the hyperprior compression model with the autoregressive regularizer; (b) the autoregressive compression model with the hyperprior regularizer.
Effect of $\alpha$ and longer training.
In Fig.(6)(a), we continue training both the unregularized and regularized hyperprior models to $4\times 10^{6}$ steps and demonstrate the effect of the regularization factor $\alpha$ on the performance. Herein, with extended training, the regularization consistently enhances optimization. The best performance, with $\alpha=1$, is achieved at the $2.5\times 10^{6}$ step, showing a BD-Rate of -1.24%. As training progresses to $4\times 10^{6}$ steps, the gain reduces to around 0.7%. For the effect of hyperparameter $\alpha$, it is shown that either the value is too large or too small will not benefit the optimization, which can be explained as follows. As discussed in Sec. 3.2, our $\alpha H(\bm{X}|\bm{\hat{X}})$ maximization regularization is essentially a bound upper of the ground-truth objective $\alpha(H(\bm{X}|\bm{\hat{X}})-H(\bm{U}|\bm{\hat{X}}))$. When $\alpha$ is large, this $H(\bm{U}|\bm{\hat{X}})$ gap is also amplified, which harms the regularization accuracy and thus the optimization performance. When $\alpha$ is small, the regularizer has little contribution to the overall loss, and the optimization is reduced to the vanilla objective.
On the alignment of the entropy models.
In our experimental setup, we emphasize the advantages of aligned entropy modeling between latent and source distributions. Herein, we extend the analysis to explore cases where the entropy models are unaligned, investigating two scenarios: a compression model equipped with (a) a stronger source entropy model and (b) a weaker source entropy model. In particular, for the source models of the hyperprior and autoregressive depicted in Fig.3(a) and Fig.3(b), respectively, we consider the autoregressive entropy model with additional spatial contexts to be a more accurate and thus “stronger” model. Thereafter, we train the hyperprior compression model with a autoregressive source regularizer and the autoregressive compression model with a hyperprior source regularizer, and the results are depicted in Fig.6(b) and Fig.6(c), respectively. The values of $\alpha$ are still set to 0.1 and 1 for the hyperprior and autoregressive regularizers, respectively. The performance of the aligned entropy model is also included for reference. From Fig.6, it can be observed that neither a stronger nor a weaker source entropy model consistently facilitates the optimization performance of the compression models. Specifically, when a stronger autoregressive regularizer is imposed on the hyperprior network, the average BD-Rate remains positive throughout training. When a weaker hyperprior regularizer is employed by the autoregressive network, the optimization is temporarily improved at the $1.5\times 10^{6}$ training step, with a BD-Rate of -0.20%. In contrast, aligning the entropy models leads to significant performance improvements.
4.3 Limitations
Our regularization method is applied exclusively during the network training stage and therefore introduces no additional inference complexity. However, since it involves training an additional source entropy model $q_{\theta}(\bm{X}|\bm{\hat{X}})$, the training complexity is inevitably increased. A straightforward yet meaningful analysis can be provided as follows. In our experiments, we demonstrate that the best regularization performance is achieved when the neural architectures of the source and latent entropy models are aligned. Assuming the complexities of the source and latent models are approximately equal, the complexity ratio can be calculated as:
$\frac{2C_{L}+C_{T}}{C_{L}+C_{T}}=1+\frac{1}{1+C_{T}/C_{L}},$ (17)
where $C_{L}$ and $C_{T}$ denote the complexity of the latent entropy model and the transform backbone, respectively. From Eqn.(17), it is indicated that the complexity is highly related to $C_{T}/C_{L}$. When the transform backbone is significantly more complex than the latent entropy model, i.e., $C_{T}/C_{L}\to\infty$, the overhead becomes less prominent. For $\lambda=0.0018$, we train $1\times 10^{5}$ steps for each model on the same machine (Nvidia RTX 4090) and compute the time ratio. As shown in Table. 2, compared to the unregularized baseline, the training time of our proposed method increases by 28.8%, 18.2%, 8.1%, 39.3% and 29.8% for the hyperprior, autoregressive, attention, ELIC and MLIC++ models, respectively. Notably, for the attention model, the overhead is the least significant. This can be attributed to the fact that the attention-based transform backbone is significantly more complex than the entropy model, as illustrated in Fig. 3(b). To further reduce training complexity, one can employ lightweight convolutional layers, such as depthwise separable convolution [56], or improve the iterative training strategy outlined in Algorithm 1.
@@ 4.3 Limitations
hyperprior autoregressive attention ELIC MLIC++
Training time 128.8% 118.2% 108.1% 139.3% 129.8%
@@ 5 Conclusion
This paper investigates a structural entropy regularizer for lossy neural image compression. Specifically, motivated by the information-theoretic analysis of compression models, we demonstrate that, in addition to minimizing latent entropy, maximizing conditional source entropy can be equally important. Subsequently, by regularizing the neural networks to simultaneously maximize the conditional source entropy during end-to-end optimization, we experimentally observe improvements in both the in-domain compact representation and the out-of-domain generalization abilities of compression models. Additionally, as the primary overhead of this regularizer comes from training an additional source entropy model, the proposed method imposes no added inference complexity, opening up a new space for regularizing neural compression networks.
@@ 6 Proof of Lemma 1
Substituting Eqn.(22) into Eqn.(23), we conclude
@@ 7 Proof of Lemma 2
Recalling the proof in Sec. 6, since both the analysis transform $\mathrm{T}_{\mathrm{A}}(\cdot)$ and synthesis transform $\mathrm{T}_{\mathrm{S}}(\cdot)$ are deterministic, the following holds:
@@
Figure 7: Details of (a) Four convolutional modules. (b) Inter attention module, (c) Intra attention module, and (d) Checkerboard attention module from MLIC++ [12].
@@ 8 Reproduced baselines
We retrain the hyperprior [4], autoregressive [8] and attention [10] models from scratch, adhering to the default implementation and training configurations of CompressAI’s [44]. Four bit-rate points, i.e., $\lambda\in\{0.0018,0.0035,0.0067,0.0130\}$ are trained with $2\times 10^{6}$ steps. The evaluation results are summarized in Table 3. On average, when compared to the pre-trained models of CompressAI’s, our reproduction models yield 0.39%, 0.41%, and -4.11% BD-Rates [57] for the hyperprior, autoregressive, and attention models, respectively. Notably, on the CLIC dataset, our reproduction models outperform their pre-trained counterparts by -1.24%, -1.56%, and -5.68% BD-Rates for the hyperprior, autoregressive, and attention models, respectively. For the attention model, our reproduction model outperforms the pre-trained model on all three evaluation datasets of Kodak, CLIC, and Tecnick by -3.32%, -3.33%, and -5.68% BD-Rates, respectively. Note that in the original attention network [10], the latent is modeled as the mixture of Gaussian with 3 clusters. In CompressAI’s default implementation, this cluster number is simplified to 1. We follow this latent implementation. The detailed rate-distortion curves for the Kodak dataset are visualized in Fig.8.
@@
9 More details on source entropy models
Herein, the module designs are identical to the latent designs [4, 8, 10, 43, 12], with only minor dimension adjustment. The details are depicted in Fig.7. For the attention modules, the depthwise separable convolution is adopted [56].